can someone please breakdown the process of elimination in algerbra???? here is an example the teacher gave me please help me break this down so I can understand how the teacher got this answer 5x+3y =_11 and 7x-2y=19 the teacher says that this is the answer ( 35/31 -172/31) how did he get this?
Answer by checkley77(12844)

5x+3y=-11 multiply by 2
7x-2y=19 multiply by 3 & add.
10x+6y=-22
21x-6y=57
-----------------
31x=35
x=35/31 ans.
5*35/31+3y=-11
175/31+3y=-11
3y=-11-175/31
3y=(-11*31-175)/35
3y=(-341-175)/31
3y=-516/31
y=(-516/31)/3
y=-516/31*1/3
y=-516/93
y=-172/31 ans.